OpenStudy (selena2345):

Determine which equation is an identity or whether it has no solution. 9x+3x-10+3(3x+x)

OpenStudy (selena2345):

Can You Please Help Me ? @satellite73

OpenStudy (anonymous):

i can't because it is not an equation it has no equal sign in it

OpenStudy (bibby):

For reference an equation says One thing is equal to the other A=B. A+7=B etc.

OpenStudy (selena2345):

oh yes there is sorry i forgot to put it @satellite73 9x+3x-10=3(3x+x)

OpenStudy (anonymous):

combine like terms on the left what is \(9x+3x\) ?

OpenStudy (selena2345):

12x @satellite73

OpenStudy (anonymous):

yes, so the left hand side is \(12x-10\) the right hand side is \(3(3x+x)=3\times 4x=12x\)

OpenStudy (anonymous):

so your equation is \[12x-10=12x\] which is not possible

OpenStudy (selena2345):

what do u mean not possible @satellite73

OpenStudy (anonymous):

is is not possible for \(12x-10\) to be equal to \(12x\)

OpenStudy (anonymous):

because one number cannot be 10 less than itself!

OpenStudy (anonymous):

oh i see what you might mean i wrote "not possible" but your answer should be "no solution"

OpenStudy (selena2345):

can u show me how u got no solution ? @satellite73

OpenStudy (e.mccormick):

= mrans equality or that they match, have the same value. So, it must be true. Is this true? 5+2=5

OpenStudy (selena2345):

no @e.mccormick

OpenStudy (e.mccormick):

By the same rule, the identity of: 5x+2=5x is not true. And 12x-10=12x is not true. That is because in an identity you can not move things across the = sign to solve for them.

OpenStudy (selena2345):

can u show me all the work ? @e.mccormick

OpenStudy (e.mccormick):

Satelite73 did. The work in an identity is to get one side to look exactly like the other. This was impossible because the result is not true. Therefore there is no solution.

OpenStudy (e.mccormick):

http://sketchtoy.com/55489118